OpenStudy (anonymous):

Find the form of the partial fraction decomposition of the following rational function. ((x-1)(x+5)(x^2+17)^2)/ (x^3(x+1)^2(x^2+8))

hartnn (hartnn):

where exactly are u stuck ??

OpenStudy (anonymous):

I don't know where to start

hartnn (hartnn):

this is very lengthy u have solved any partial fractions problem earlier ?

OpenStudy (anonymous):

no this is the first one, and its due within an hour :(

hartnn (hartnn):

do u have options ?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

it says i do not have to determine the values for a 1 or a 2

hartnn (hartnn):

choices ... something to choose from.

hartnn (hartnn):

where does a1 or a2 come from ?

OpenStudy (anonymous):

Find the form of the partial fraction decomposition of the following rational function. You do NOT need to determine the values of the constants A1,A2, . . . (or A,B, . . .).

OpenStudy (anonymous):

that is the entire question

hartnn (hartnn):

okk, then its not much lengthy u have denominator as

hartnn (hartnn):

\(x^3(x+1)^2(x^2+8)\) right ?

OpenStudy (anonymous):

yes that is denominator

hartnn (hartnn):

if only x^3 then u write \(A1/x+A2/x^2+A3/x\) ok ?

OpenStudy (anonymous):

ok, where should i write that at

hartnn (hartnn):

wait, for (x+1)^2 u have \(A4/(x+1)^2+A5/(x+1)\) and for x^2+8 u have \(\frac{A6x+A7}{x^2+8}\) now just add all of them...

OpenStudy (anonymous):

so what would it look like in the end?

hartnn (hartnn):

\(\huge\frac{A_1}{x}+\frac{A_2}{x^2}+\frac{A_3}{x^3}+\frac{A_4}{(x+1)^2}+\frac{A_5}{x+1}+\frac{A_6x+A_7}{x^2+8}\)

OpenStudy (anonymous):

is this the work?

hartnn (hartnn):

yup. your final answer.

OpenStudy (anonymous):

but how can it be that easy though, all you did was just spread out the values????

hartnn (hartnn):

its easy because you don't have to find the values of constants, just know how to decompose

OpenStudy (anonymous):

so it was just a one step process then?

hartnn (hartnn):

yes, if u need to show work, show it separately for each denominator , then add.

OpenStudy (anonymous):

can you just show me an example for the first term then on how to do that?

hartnn (hartnn):

like how to find values of constants ??

OpenStudy (anonymous):

the showing the work part

hartnn (hartnn):

see my 1st comment after your comment 'yes that is denominator' thats where i have shown it, in one line....A1/..........

OpenStudy (anonymous):

thanks man

hartnn (hartnn):

welcome :)